here is the problem: Using the rule of 72, calculate how quickly $100 will double to $200 at interest rates of 2%, 4%, 6%, 8% and 10%.

Using the rule of 72, calculate how quickly $100 will double to $200 at interest rates of 2%, 4%, 6%, 8% and 10%.
Answer #1

how quickly $100 will double to $200 at interest rates of 2%

=72/2

=36 years

how quickly $100 will double to $200 at interest rates of 4%

=72/4

=18 years

how quickly $100 will double to $200 at interest rates of 6%

=72/6

=12 years

how quickly $100 will double to $200 at interest rates of 8%

=72/8

=9 years

how quickly $100 will double to $200 at interest rates of 10%

=72/10

=7.2 years


answered by: AceMoon